OpenStudy (anonymous):

Can someone please help me solve this: In 1920, the record for a certain race was 46.7 sec. In 1930 it was 46.5 sec. Let R(t)=the record in the race and t=the number of years since 1920. Find a linear function that fits the data(Round to the nearest hundredth)

OpenStudy (anonymous):

easy cake. change in R(t) for every 10 increments of t is -.2 so -.2/10=-.02 R(t)=-.02t+46.7

OpenStudy (anonymous):

yes this was easy

OpenStudy (anonymous):

ok now i need to know how to figure out the predicted record for 2003

OpenStudy (anonymous):

and 2006

OpenStudy (anonymous):

I think the correct answer would be 45.04 is that right?

OpenStudy (anonymous):

hmmm? sorry was away. 2003-1920=83 2006-1920=86 plug those in to t

OpenStudy (anonymous):

first one's right second one's 44.98

OpenStudy (anonymous):

ok that's what I came up with too

OpenStudy (anonymous):

nice medal me =D

OpenStudy (anonymous):

ok now my next task is to figure out when the recod will be 44.76 sec

OpenStudy (anonymous):

so R(t)=44.76 44.76=-.02t+46.7 solve for t